# Design of A-Star based heuristic algorithm for efficient interdiction in multi-Layer networks

### v2

@@ 2 Problem description
We consider an escape interdiction problem involving multiple defenders, denoted by $\overline{D}=\{d_{r}\mid r\in R\}$, and a single attacker $\overline{A}$. The total number of defenders is $R$, and the set of all defenders is represented by $\overline{D}$. The attacker has a finite set of possible actions, denoted by $\acute{A}$, which is provided as input. Our objective is to compute the optimal strategy for the defenders using the proposed algorithm. The defenders act jointly, and a complete defender strategy comprises the strategies of all individual defenders.
@@
The utility of the defender is computed using the probability distribution $y$ over $\acute{A}$, as shown in Eq. (1). Considering the developed defender strategy $x$, the best defender utility is defined as:
$U_{d}(x,y)=\sum\limits_{A\in\acute{A}}U_{d}(x,A)\bullet y_{A}$ (1)
@@ 3 Proposed A-Star algorithm on multi-layer networks
       for each state $(n,t_{a})$ in the strategy do
             Assign mixed probability $P_{\text{mix}}$ to node $n$ in graph $G_{t_{a}}$ where $P_{mix}$ is the mixed probability of the current attacker strategy ;
@@
for $(i=0;i<t_{\text{max}}-1;i++)$ do
       for $(j=i;j<t_{\text{max}};j++)$ do
             if edge exists between same node in $G_{j}$ and $G_{j+1}$ then
                   Update the exact value of the node at $G_{j+1}$ = exact value of the node at $G_{j+1}$ + exact value of the node at $G_{j}$ ;
@@
for $(k=t_{\text{max}}-2;k\geq 0;k--)$ do
       if path exists from node $n$ in $G_{k}$ to same node $n$ in $G_{t_{\text{max}}-1}$ then
             $h(n)$ at $G_{k}$ = ($g(n)$ at $G_{t_{\text{max}}-1}$) - ($g(n)$ at $G_{k}$);
             for each incoming edge to this node $n$ in $G_{k}$ do
                   Propagate $h(n)$ to the source node of the edge;

### v3

@@ 2 Problem description
We consider an escape interdiction problem involving multiple defenders, denoted by $\overline{D}=\{d_{r}\mid r\in R\}$, and a single attacker $\overline{A}$. The total number of defenders is $m$, and the set of all defenders is represented by $\overline{D}$. Here, $r\in R=\{1,\dots,m\}$. The defender’s pure strategy is denoted by $S$. A pure strategy for the defender consists of $m$ schedules, i.e., $S=\{S^{r}:r\in R\}$. The attacker has a finite set of possible actions, denoted by $\acute{A}$, which is provided as input. Our objective is to compute the optimal strategy for the defenders using the proposed algorithm. The defenders act jointly, and a complete defender strategy comprises the strategies of all individual defenders.
@@
The utility of the defender is computed using the probability distribution $y$ over $\acute{A}$, as shown in Eq. (1). Considering the developed defender strategy $S$, the best defender utility is defined as:
$U_{d}(S,y)=\sum\limits_{A\in\acute{A}}U_{d}(S,A)\bullet y_{A}$ (1)
@@ 3 Proposed A-Star algorithm on multi-layer networks
       for each state $(n,t_{a})$ in the current attacker strategy do
             Exact cost $g(n)$ of node $n$ in graph $G_{t_{a}}$ = $g(n)$ + $P_{\text{mix}}$, where $P_{mix}$ is the mixed probability of the current attacker strategy ;
@@
Update exact cost of GOAL nodes:
$min\_t\_index$ = $M$ where $M$ is large integer value ;
for each attacker strategy do
       for each state $(n,t_{a})$ in the current attacker strategy do
             if $n$ == GOAL node then
                   if $min\_t\_index>t_{a}$ then
                         $min\_t\_index$ = $t_{a}$;
@@
             end if
[…]
                   Update the exact value of the GOAL node at $G_{i+1}$ = exact value of the GOAL node at $G_{i+1}$ + exact value of the GOAL node at $G_{i}$ ;
@@
end for
Update exact cost of all non-interdicted nodes:
if Exact cost $g(n)$ of any node == 0 then
       $g(n)$ of that node = $\infty$;
end if
@@
for $(k=t_{\text{max}}-2;k\geq min\_t\_index;k--)$ do
       Heuristic cost $h(n)$ of GOAL node at $G_{k}$ = exact cost ($g(n)$ of GOAL node at $G_{t_{\text{max}}-1}$) - exact cost of GOAL node ($g(n)$ at $G_{k}$);
       for each incoming edge to the GOAL node $n$ in $G_{k}$ do
             Propagate $h(n)$ as the heuristic cost of the source node of the edge and repeat this same procedure for all incoming edges considering this source node as destination node;